OpenStudy (anonymous):

P^4- 81= ( p^2-9)(p^2+9) = (p+3)(p-3)(p^2+9); Why can't you further factor (P^2+9)? By factoring by grouping P^2+0P+9, I get (P+3)(P-3), So my complete answer is (P+3)(P-3)(P+3)(P-3).

OpenStudy (phi):

notice that (p+3)(p-3) = p^2 - 9 which is different from p^2 + 9 in other words, p^2 + 9 does not factor

OpenStudy (anonymous):

I see it now. I couldn't for the life of me figure it out. Thank you so much!!

OpenStudy (anonymous):

if we consider p^2+9 as quadratic polynomial then discriminent=0^2-4*1*9=-36 which is negative so can not be factored for real numbers.

OpenStudy (anonymous):

I haven't gotten into quadratic equations just yet, it's coming up in the next section though. Wish me luck!

OpenStudy (phi):

luck doesn't work with math (unfortunately). But hard work and concentration do.

OpenStudy (anonymous):

I know, I've been dedicated to my math class and am among the top students in my class. :)

OpenStudy (anonymous):

Thanks again, Phi! :D